OpenStudy (anonymous):

Simplify completely quantity 5 x squared plus 9 x minus 2 over quantity x squared plus 12 x plus 20 times quantity x squared plus 17 x plus 70 over quantity 15 x minus 3.

OpenStudy (anonymous):

Is this correct? \[\frac{ 5x^2 + 9x - 2 }{ x^2 + 12x + 20 } \times \frac{ x^2 + 17x + 70 }{ 15x - 3 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Try factorising each part of the expression, you'll find some factors on the top and bottom of the fraction that cancel out.

OpenStudy (anonymous):

Ok I did but I think im doing something wrong

OpenStudy (anonymous):

What did you get when you factorised it all?

OpenStudy (anonymous):

\[5x2+9x-2 \] \[(x+2) (x+10)\]

OpenStudy (anonymous):

\[(x+10) ( x+7)\] \[3(5x)\]

OpenStudy (anonymous):

The first one can be factorised, so you get \[5x^2 + 9x - 2 = (5x-1)(x+2)\] and you forgot a term on the denominator of the 2nd bit, you should get \[3(5x-1)\] instead of \[3(5x)\] When you change that, you should get a bunch of cancellation :)

OpenStudy (anonymous):

so my answer should be (x+7) over 3

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

This is correct! The answer was A) (x+7)/3